What quantity of 79% hydrochloric acid solution must be mixed with a 16% hydrochloric acid solution to produce 3729.6 mL of 21% solution
Found 2 solutions by mananth, addingup:
Answer by mananth(16946)

percent ---------------- quantity
Acid type I 79.00% ---------------- x liters
acid type II 16.00% ------ 3729.6 - x liters
Mixture 21.00% ---------------- 3729.6
3729.6
79.00% x + 16.00% ( 3729.6 - x ) = 21.00% * 3729.6
79 x + 16 ( 3729.6 - x ) = 78321.6
79 x + 59673.6 - 16 x = 78321.6
79 x - 16 x = 78321.6 - -59673.6
63 x = 18648
/ 63
x = 296 ml 79.00% Acid type I
3433.6 ml 16.00% acid type II

m.ananth@hotmail.ca

Answer by addingup(3677)

0.79x+0.16(3729.6-x)= 3729.6(0.21)
0.79x+596.74-0.16x = 783.22
0.63x = 186.48
x = 296mL at 79% and the rest at 16%
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
Check:
296*0.79 = 233.84
3729.6-296 = 3433.60
3433.60*0.16 = 549.38
Add them up: 233.84+549.38 = 783.22 Correct because the problem says that you want 3729.6 at 21% and 3729.6*0.21= 783.22